A parallel-plate capacitor has 2.00 cm2 plates that are separated by 5.00 mm with air between them. If a 12.0 V battery is connected to this capacitor, how much energy does it store?
Answer #1
energy stored
E = (1/2)C * V^2
C = (εoA/d)
εo= 8.85 * 10^-12 C^2/Nm^2
A = 2.00 cm^2 = 2.00 * (10^-2)^2 m^2 = 2.00 * 10^-4 m^2
d = 5.00 mm = 5.00 * 10^-3 m
V = 12.0 V
E=(1/2)* [ 8.85*10^(-12) * 2.00*10^(-4)/5*10^-3]*12^2 = 2.5488 x 10^(-11)
Check the calcs again! It should be ok but go over it just in case.
Please, rate! Thank you.